OpenStudy (anonymous):

xy=4 whats dx/dt when x=1 and dy/dt=-6

OpenStudy (solomonzelman):

rephrase, please!

OpenStudy (solomonzelman):

What are the equations? and what are you looking for?

OpenStudy (solomonzelman):

He left! Why?

OpenStudy (anonymous):

im looking for dx/dt

OpenStudy (solomonzelman):

Do you guys get the question?

OpenStudy (solomonzelman):

when xy=4 and x=1?

OpenStudy (anonymous):

its 3/2

OpenStudy (anonymous):

can you help me with the differentiating part

OpenStudy (solomonzelman):

xy=4 x=1 dy/dt=-6

OpenStudy (anonymous):

use d(xy)=x.y' + x'y=d(4) and xy=4 as x=1 y=4

OpenStudy (solomonzelman):

So... xy=4, x=1 plug in 1 for x, (1)4=y, y=4 d4/dt=6 d s cancel

OpenStudy (solomonzelman):

@chels.408 do you need more help with understanding this?

OpenStudy (solomonzelman):

You have: xy=4 x=1 dy/dt=-6 Looking for: dx/dt

OpenStudy (anonymous):

yeah but its a related rates problem

OpenStudy (solomonzelman):

x=1, plug that into xy=4 (1)y=4 so y=4 NOW: dy/dt=-6 plug 4 for y into the above equation. 4d/dt=-6 cancel out Dees 4/t=-6 t=-3/2 Ur looking for: dx/dt So plug in 1 for x and -3/2 for t